# The impact of cosmic filaments on the abundance of satellite galaxies

### v1

@@ 4.1 Dark matter-traced filaments
4.2 Tracer effects on satellite luminosity functions
Once the dark matter-traced filament spines are identified, we follow the same environmental definitions outlined in Section 2.3 and apply the analysis described in Section 3, as done for the galaxy-traced case. The numbers of central galaxies classified into filament and field environments across different magnitude bins are given in the two rightmost columns of Table 1. Compared to the galaxy-traced case, a larger number of central galaxies are assigned to the filament environment, reflecting the fact that dark matter-traced filaments can capture more low-density ridges and thus include the galaxies surrounding these structures. The resulting environmental dependence of satellite LFs is summarized in Figure 6 and Table 2.
@@ 4.2 Tracer effects on satellite luminosity functions
4.3 How do galaxy tracers enhance the satellite abundance in filaments?
@@
The effects of tracer choice on satellite LFs, outlined in Section 4.2, are closely related to the properties of filament-finding algorithms. Algorithms such as DisPerSE  construct the cosmic web based on the spatial distribution of tracer particles. They estimate a density field from the tracer distribution and identify filaments as ridge-like structures connecting local density maxima. As a result, regions with more densely clustered tracers yield higher estimated densities and are more likely to be identified as part of the filamentary network. Consider two halos with similar virial masses residing in comparable large-scale environments – one hosting more satellites (satellite-rich) and the other fewer (satellite-poor), although their actual surrounding dark matter densities may be similar, the satellite-rich halo contributes more to the galaxy-traced density field and is therefore more likely to be classified as residing within a galaxy-traced filament.
@@
To illustrate this intuitively, we construct controlled central galaxy samples – satellite-rich and satellite-poor – based on the filament galaxy samples from both the galaxy-traced and dark matter-traced cases described earlier. We adopt the same three central galaxy magnitude bins used previously: $M_{r,{\rm cen}}\sim-22$, 21-21, and 20-20. In each magnitude bin, satellite-rich central galaxies are defined as those hosting at least one satellite with $M_{\rm diff}<3$, while satellite-poor central galaxies host none. Both central galaxy samples are then further resampled to match the $M_{r,{\rm cen}}$ and virial mass $M_{\mathrm{200c}}$ distributions, ensuring a fair comparison.
@@
For the galaxy-traced filaments, the number ratios between satellite-rich and satellite-poor central galaxies are $N_{\rm rich}/N_{\rm poor}=$ 1.28, 1.37, and 1.25 in the $M_{r,{\rm cen}}\sim-22$, 21-21, and 20-20 bins, respectively. In contrast, these ratios are $N_{\rm rich}/N_{\rm poor}=$ 0.97, 0.96, and 1.05 in the dark matter-traced filaments (see markers in Figure 7). This clearly shows that, compared to the dark matter-traced case, the galaxy-traced filaments tend to pass through central galaxies hosting more satellites. This explains the more pronounced environmental differences in satellite LFs observed for the galaxy-traced case in previous subsections. Additionally, the fact that the ratio $N_{\rm rich}/N_{\rm poor}$ is quite close to $1$ in the dark matter-traced case further suggests that dark matter density is a less biased tracer of the underlying filamentary network.
@@ 5 Summary
We acknowledge support from the National Key Program for Science and Technology Research and Development of China (2023YFB3002500), the National Natural Science Foundation of China (NSFC; No. 12588202) and the China Manned Space Project (CMS-CSST-2025-A03). SL acknowledges the support by the NSFC grant (No. 12473015). LW acknowledges support from the National SKA Program of China (No.2022SKA0110201), and the National Key Research and Development Program of China (No.2023YFB3002501).

### v2

@@ IV.1 Dark matter-traced filaments
IV.2 Tracer effects on satellite luminosity functions
Once the dark matter-traced filament spines are identified, we follow the same environmental definitions outlined in Section II.3 and apply the analysis described in Section III, as done for the galaxy-traced case. The numbers of central galaxies classified into filament and field environments across different magnitude bins are given in the two rightmost columns of Table 1. We note that the total number of central galaxies in the filament plus field environments is identical to that in the galaxy-traced case owing to our environment definition (see Section II.3). Compared to the galaxy-traced case, a larger number of central galaxies are assigned to the filament environment, reflecting the fact that dark matter-traced filaments can capture more low-density ridges and thus include the galaxies surrounding these structures. The resulting environmental dependence of satellite LFs is summarized in Figure 6 and Table 2.
@@ IV.2 Tracer effects on satellite luminosity functions
IV.3 How do galaxy tracers enhance the satellite abundance in filaments?
@@
The effects of tracer choice on satellite LFs, outlined in Section IV.2, are closely related to the properties of filament-finding algorithms. Algorithms such as DisPerSE  construct the cosmic web based on the spatial distribution of tracer particles. They estimate a density field from the tracer distribution and identify filaments as ridge-like structures connecting local density maxima. As a result, regions with more densely clustered tracers yield higher estimated densities and are more likely to be identified as part of the filamentary network. Consider two halos with similar virial masses residing in comparable large-scale environments – one hosting more satellites (satellite-rich) and the other fewer (satellite-poor), although their actual surrounding dark matter densities may be similar, the satellite-rich halo contributes more to the galaxy-traced density field and is therefore more likely to be classified as residing within a galaxy-traced filament.
@@
To illustrate this intuitively, we construct controlled central galaxy samples – satellite-rich and satellite-poor – from the total central galaxy sample after excluding those in the knot environment (see Section II). We adopt the same three central galaxy magnitude bins used previously: $M_{r,{\rm cen}}\sim-22$, 21-21, and 20-20. In each magnitude bin, satellite-rich central galaxies are defined as those hosting at least one satellite with $M_{\rm diff}<3$, while satellite-poor central galaxies host none. [^3] We also tested higher $M_{\rm diff}$ thresholds to include fainter satellites (e.g. $M_{\rm diff}<6$) and considered defining satellite-rich centrals as hosting at least one or two satellites. In all cases, the main conclusions remain consistent. We then further resample the satellite-rich and satellite-poor central galaxy samples to match their $M_{r,{\rm cen}}$ and virial mass $M_{\rm 200c}$ distributions, ensuring a fair comparison. Finally, we examine the numbers of satellite-rich and satellite-poor central galaxies in this controlled sample that reside in the identified filaments, denoted $N_{\rm rich}$ and $N_{\rm poor}$, respectively.
@@
For the galaxy-traced filaments, the number ratios between satellite-rich and satellite-poor central galaxies are $N_{\rm rich}/N_{\rm poor}=$ 1.28, 1.37, and 1.25 in the $M_{r,{\rm cen}}\sim-22$, 21-21, and 20-20 bins, respectively. In contrast, these ratios are $N_{\rm rich}/N_{\rm poor}=$ 0.97, 0.96, and 1.05 in the dark matter-traced filaments (see markers in Figure 7).  [^4] To ensure that difference in the magnitude distributions of central galaxies (the combined satellite-rich and satellite-poor populations) between galaxy-traced and dark matter-traced filaments does not bias the results, we performed a supplemental resampling test. After matching both samples in $M_{r,{\rm cen}}$ and $M_{\rm 200c}$, the resulting $N_{\rm rich}/N_{\rm poor}$ ratios remained consistent with those reported in the main text. This clearly shows that, compared to the dark matter-traced case, the galaxy-traced filaments tend to pass through central galaxies hosting more satellites. This explains the more pronounced environmental differences in satellite LFs observed for the galaxy-traced case in previous subsections. Additionally, the fact that the ratio $N_{\rm rich}/N_{\rm poor}$ is quite close to $1$ in the dark matter-traced case further suggests that dark matter density is a less biased tracer of the underlying filamentary network.
@@ V Summary
We thank the anonymous referee for their constructive comments, which helped improve this manuscript. We acknowledge support from the National Natural Science Foundation of China (NSFC) grant (No. 12588202), the National Key Program for Science and Technology Research and Development of China (2023YFB3002500), and the China Manned Space Project (CMS-CSST-2025-A03). SL acknowledges the support by the NSFC grant (No. 12473015). LW acknowledges support from the National SKA Program of China (No.2022SKA0110201), and the National Key Research and Development Program of China (No.2023YFB3002501).
Appendix A Effects of filament persistence thresholds
Filament identification with DisPerSE depends on the choice of the persistence ratio, expressed in units of $\sigma$. In the main text, we adopt a fiducial threshold of $2\sigma$. Higher thresholds select fewer filaments and filter out unreliable or spurious structures, which results in shorter total filament lengths. Conversely, lower thresholds include more filaments and lead to longer total lengths. As a result, both the number of galaxies classified as in filaments and the corresponding satellite LFs could be affected by the choice of persistence threshold.
Here we test the robustness of our conclusions using persistence thresholds of $1\sigma$, $4\sigma$, and $6\sigma$. The total filament lengths, $L_{\rm tot}$, at $1\sigma$, $4\sigma$, and $6\sigma$ are approximately $1.65$, $0.65$, and $0.26$ times the fiducial $L_{\rm tot,fiducial}$ at $2\sigma$, respectively. The persistence ratios for the dark matter–traced filaments are chosen to match the total filament lengths, as described in Section IV.1.
Figures 8, 9, and 10 show the environmental differences in the satellite LFs for different persistence thresholds. Relative to Figures 4 and 6 (the $2\sigma$ results), the satellite LFs and the ratios between filament and field environments remain broadly similar. A more noticeable deviation appears at $M_{r,{\rm cen}}\sim-22$ (original sample, red curve) in the $6\sigma$ case, where many filaments become classified as field regions, causing the environmental ratio to drop. In all thresholds considered, resampling in $M_{r,{\rm cen}}$ and $M_{\rm 200c}$ suppresses the environmental differences, and using dark matter as the filament tracer further reduces them.
For a quantitative comparison, we summarize the average ratios in Table 3, analogous to Table 2. We find that the average satellite abundance ratio correlates with the persistence threshold: lower persistence (more filaments and longer total length) corresponds to a higher ratio, and vice versa. However, the fractional reduction from resampling remains similar across thresholds. Resampling in $M_{r,{\rm cen}}$ reduces the ratio by $\sim 0-30\%$ (16–84th percentile), with a median reduction of $\sim 15\%$. Resampling in both $M_{r,{\rm cen}}$ and $M_{\rm 200c}$ reduces the ratio by $\sim 50-90\%$ (16–84th percentile), with a median of $\sim 75\%$. For fixed total filament length, dark matter–traced filaments consistently show lower satellite abundance ratios than galaxy-traced filaments.
In summary, varying the persistence ratio from $1\sigma$ to $6\sigma$ changes the total filament length and correspondingly affects the satellite LFs and abundance ratios. Nevertheless, the main conclusions regarding environmental effects and tracer dependence remain robust.
Figure 8: Environmental differences in satellite LFs, analogous to Figures 4 and 6, but using filaments identified with a persistence threshold of $1\sigma$. Left: Galaxy-traced filaments without resampling. Middle: Galaxy-traced filaments after resampling in $M_{r,{\rm cen}}$ and $M_{\rm 200c}$. Right: Dark matter–traced filaments ($4.4\sigma$) after the same resampling. The satellite LFs and their filament-to-field ratios are quantitatively different compared to the fiducial $2\sigma$ case, but the behaviors are qualitatively similar, especially the effects of halo mass and tracer choice are consistent with the results discussed in the main text.
Figure 9: Same as Figure. 8, but for filaments identifies with a higher persistence threshold of $4\sigma$.
Figure 10: Same as Figure. 8, but for filaments identifies with a higher persistence threshold of $6\sigma$.
Table 3: Average satellite abundance ratios between filament and field central galaxies, analogous to Table 2, but for filaments identified with different persistence thresholds ($1\sigma$, $4\sigma$, and $6\sigma$). Note that the original ratio in the last row does not exceed unity, so the calculation of the fractional reduction is not applicable and is therefore omitted.
Filament tracers Bins $\phi_{\rm sat,filament}/\phi_{\rm sat,field}$
Original Resampling $M_{r,{\rm cen}}$ Resampling $M_{r,{\rm cen}}$ and $M_{\rm 200c}$
Galaxies ($1\sigma$) $M_{r,{\rm cen}}\sim-22$ $3.62$ $3.19~(17\%\downarrow)$ $1.60~(77\%\downarrow)$
$M_{r,{\rm cen}}\sim-21$ $2.73$ $2.59~(8\%\downarrow)$ $1.44~(74\%\downarrow)$
$M_{r,{\rm cen}}\sim-20$ $1.89$ $1.89~(0\%\downarrow)$ $1.40~(56\%\downarrow)$
Dark matter density ($4.4\sigma$) $M_{r,{\rm cen}}\sim-22$ $3.26$ $2.42~(37\%\downarrow)$ $1.26~(89\%\downarrow)$
[…]
$M_{r,{\rm cen}}\sim-20$ $1.64$ $1.54~(15\%\downarrow)$ $1.23~(63\%\downarrow)$
Galaxies ($4\sigma$) $M_{r,{\rm cen}}\sim-22$ $3.10$ $2.76~(16\%\downarrow)$ $1.50~(76\%\downarrow)$
$M_{r,{\rm cen}}\sim-21$ $2.59$ $2.46~(9\%\downarrow)$ $1.29~(82\%\downarrow)$
$M_{r,{\rm cen}}\sim-20$ $1.87$ $1.99~(14\%\uparrow)$ $1.44~(50\%\downarrow)$
Dark matter density ($5.6\sigma$) $M_{r,{\rm cen}}\sim-22$ $1.61$ $1.35~(42\%\downarrow)$ $1.09~(85\%\downarrow)$
$M_{r,{\rm cen}}\sim-21$ $1.36$ $1.29~(19\%\downarrow)$ $1.01~(98\%\downarrow)$
[…]
Galaxies ($6\sigma$) $M_{r,{\rm cen}}\sim-22$ $1.86$ $1.47~(46\%\downarrow)$ $1.22~(75\%\downarrow)$
$M_{r,{\rm cen}}\sim-21$ $2.65$ $2.43~(14\%\downarrow)$ $1.32~(81\%\downarrow)$
$M_{r,{\rm cen}}\sim-20$ $2.20$ $2.47~(22\%\uparrow)$ $1.64~(46\%\downarrow)$
Dark matter density ($6.5\sigma$) $M_{r,{\rm cen}}\sim-22$ $1.37$ $1.28~(26\%\downarrow)$ $1.00~(100\%\downarrow)$
$M_{r,{\rm cen}}\sim-21$ $1.33$ $1.30~(9\%\downarrow)$ $1.02~(95\%\downarrow)$
$M_{r,{\rm cen}}\sim-20$ $0.89$ 0.89()0.89~(-) 1.08()1.08~(-)